Consider a poison distribution with a mean ¡=3 then P( X<2) is
1. 0.2240
2. 0.1992
3. 0.4232
4. 0.1494
5. 0.4223
1
Expert's answer
2020-11-09T19:37:50-0500

"P( X<2)=P( X=0)+P( X=1)\\
=\frac{e^{-3}3^{0}}{0!}+\frac{e^{-3}3^{1}}{1!}\\
\approx 0.1992"